You walk 100 yards due south, then 120 yards due west, and then 20 yards due north. How far are you from your starting point?
Answer by richwmiller(17219)

You really should draw a picture of this to understand the problem.
Use the top of the page as North the bottom as South, the left side as West and the right side as East
...............................A
...............................|
.....................B.....C|
.....................|........|
.....................|___|
draw a line from B to C to finish the rectangle
a=BC is 120 yards
b=AC 100-20=80 yards
c=AB is the hypotenuse of the triangle ABC
a^2+b^2=c^2
120^2+80^2=c^2
c=40sqrt(13)
c=144.22 yards